OpenStudy (anonymous):

How do I prove this identity? (sin x +cos x)2 = 1 + 2 sin x cos x

OpenStudy (xapproachesinfinity):

i suppose that \[\large \rm (\sin x+\cos x )^2=1+2\sin x \cos x\]

OpenStudy (xapproachesinfinity):

just expand the left hand side

OpenStudy (anonymous):

I don't know how to expand it

OpenStudy (xapproachesinfinity):

\[\large (\sin x+ \cos x)^2=(\sin x +\cos x)(\sin x+\cos x)\]

OpenStudy (xapproachesinfinity):

well if you don't know that is a big problem

OpenStudy (xapproachesinfinity):

do you know how to expand this binomial \[(a+b)^2\]

OpenStudy (anonymous):

I was in the hospital for a month so this is a pain to catch up on and no

OpenStudy (xapproachesinfinity):

even if you are in hospital does not mean you have and excuse my dear expanding a binomial is something your learn from basic algebra not even this

OpenStudy (xapproachesinfinity):

binomial expansion for (a+b)^2 is \[\large \rm (a+b)^2=(a+b)(a+b)=a^2+ab+b^2\]

OpenStudy (xapproachesinfinity):

now use the same method to expand \[\large \rm (\sin x+\cos x)^2\]

OpenStudy (welshfella):

let sinx = s and cos x = c (s + c)^2 = s^2 + c^2 + 2 sc but s^2 + c^2 = 1 so right side = 1 + 2sc 1 + 2sin x cos x

OpenStudy (xapproachesinfinity):

you shouldn't give answer that @welshfella the poster needs to work the rest himself/herself to learn

OpenStudy (welshfella):

yes ur right